User name dohrmann
Log entry time 05:24:56 on January28,2001
Entry number 55932
keyword=HV for left hrs s2.adc_r_c2 plugged in
We made a controlled access to find that the HV cable for the left HRS channel
s2.adc_r_c2 wasn't properly connected. We fixed that (thanks Joerg!),
1799-1804 were affected by this problem. Fixed from 1805 on.